Raunak took 5 moles of carbon atoms in a container arid Krish also took 5 moles of sodium atoms in another container of same weight.
(a) Whose container is heavier?
(b) Whose container has more number of atoms?

Grade:12th pass

1 Answers

Pawan Prajapati
2 years ago
(a) Mass of sodium atoms carried by Krish = (5 × 23) g = 115 g While mass of carbon atom carried by Raunak = (5 × 12) g = 60 g Thus, Krish’s container is heavy. (b) Both the bags have same number of atoms as they have same number of moles of atoms.